### v1

# A Universal Knowledge Embedded Contrastive Learning Framework for Hyperspectral Image Classification

Abstract
Hyperspectral image (HSI) classification techniques have been intensively studied and a variety of models have been developed. However, these HSI classification models are confined to pocket models and unrealistic ways of datasets partitioning. The former limits the generalization performance of the model and the latter is partitioned leads to inflated model evaluation metrics, which results in plummeting model performance in the real world. Therefore, we propose a universal knowledge embedded contrastive learning framework (KnowCL) for supervised, unsupervised, and semisupervised HSI classification, which largely closes the gap of HSI classification models between pocket models and standard vision backbones. We present a new HSI processing pipeline in conjunction with a range of data transformation and augmentation techniques that provide diverse data representations and realistic data partitioning. The proposed framework based on this pipeline is compatible with all kinds of backbones and can fully exploit labeled and unlabeled samples with expected training time. Furthermore, we design a new loss function, which can adaptively fuse the supervised loss and unsupervised loss, enhancing the learning performance. This proposed new classification paradigm shows great potentials in exploring for HSI classification technology. The code can be accessed at https://github.com/quanweiliu/KnowCL.
Index Terms:
Contrastive learning (CL), hyperspectral image (HSI) classification, semi-supervised learning, vision transformer (ViT).
publicationid: pubid:
I Introduction
The motion of electrons within matter produces electromagnetic radiation, which results in the formation of an optical spectrum. Objects typically exhibit distinct spectra, and hyperspectral imaging techniques can precisely capture these spectral variations.[1]. As a result, hyperspectral images (HSIs) have powerful recognition capability and are applied to agriculture, military, energy, and cultural relics, which makes pixel-level HSI classification a hot research topic [2, 3]. Roughly, these methods are divided into two major categories: traditional machine learning methods and deep learning methods[4]. These data-driven methods can extract different levels of features in line with the labelling of the samples. However, labelling of HSIs is expert-dependent and time-consuming, which leads to the scarcity of labeled HSIs, resulting in the current HSI classification techniques stopping at pocket models and unrealistic ways of data partitioning[5].
To fully exploit the information contained in HSIs, previous work focus on machine learning models that first extract a series of features of HSIs, such as spectral features, texture features, morphological semantic features, etc. Then the extracted feaures are fed into classifiers such as $k$-nearest neighbor classifier (KNN) to obtain classification results[6, 7, 8]. However, early machine learning methods relied on handcrafted shallow features and empirical assumptions that seriously hinder the model building efficiency and generalization performance.
Figure 1: Performance of KnowCL on DFC2018 using ViT and ResNet architecture, compared to other unsupervised and supervised baselines.
Deep learning can extract feature representations end-to-end, which was applied in HSI classification in recent years [9]. Depending on whether labels are involved in feature extraction or not, deep neural networks (DNNs) are categorized as supervised learning, unsupervised learning, and semi-supervised learning [10, 11, 12]. Minimizing the difference between the labels of the training sample and the ground truth is the learning criterion for supervised learning that often requires a large number of samples with annotations to train the model [13]. However, large-scale data labeling, especially for HSIs, is difficult to achieve [5]. Unsupervised learning, on the other hand, does not require the involvement of ground truth and it can learn discriminative feature representations from large-scale unlabeled data. Nevertheless, the way training a large amount of unlabeled hyperspectral data will cost a lot of resources. The obtained task-agnostic feature representations often struggle to outperform models dedicated to supervised learning in specific downstream tasks [14, 15]. For HSI classification tasks, we need to collect the semantic labels of samples before we predict the categories of other regions by the model. Unsupervised learning generally uses a model of unsupervised pre-training plus supervised fine-tuning, which does not fully utilize the precious label information [16].
To address the above-mentioned problems, semi-supervised learning methods that can handle labeled and unlabeled data simultaneously have been proposed. Semi-supervised deep learning is mainly implemented with the help of complex data processing and classical neural network models [17, 18, 19]. Traditional semi-supervised learning algorithms are acquired to feed the entire image into the network. Although global features can be adequately extracted, their memory footprint and computational complexity grow exponentially with the increase of data size [20]. Especially for HSIs consisting of tens of thousands of bands, these methods can only validate the model on some small datasets and cannot be applied to relatively large-size HSIs [18]. More importantly, due to the limited availability of benchmark data and the high cost of labeling, the number of parameters of the aforementioned HSI classification models has been much smaller than that of the conventional image classification models, as shown in Fig. 1. The gap has been limiting the development of HSI classification technology.
In addition to the learning strategies, another critical step is the evaluation of HSI classification methods, such as benchmark datasets, sampling strategies and metrics, where sampling strategies have become one of the most urgent methods to be revised [21]. A common sampling setting in the HSI classification is to randomly select a fixed number or a fixed proportion of labeled samples from an HSI as a training set and the rest of the samples in this image as a test set [22]. These selected samples are randomly distributed throughout the image. The spatial autocorrelation of the land cover makes similar features share the same label with a high probability [23]. Consequently, the random sampling based on pixel-by-pixel is very inefficient and impractical. In a real environment, most regions rather than pixels are usually identified as ground truth by visual interpretation or field survey. On the other hand, with the development of HSI classification based on CNNs, spectral-spatial classification becomes a concise and efficient method. However, the classification accuracy is appreciably high due to the spatial autocorrelation in the test data set. A considerable number of work has achieved nearly 100% correctness in conditions where only a small number of training samples are used [24, 25, 26, 27]. Cao et. al. [28] conducted a preliminary performance comparison of HSI classification based on the overlapped and nonoverlapped sampling strategy, where the experimental results also showed that the classification accuracy is significantly reduction using the nonoverlapped sampling strategy.
Figure 2: The KnowCL framework for HSI classification. The framework is fed with the labeled samples patch $x$, and unlabeled samples patch $\hat{x}$ in supervised training and unsupervised training, respectively. For semi-supervised training, $x$ and $\hat{x}$ are fed jointly into the network.
Given the problems with modeling and evaluation, in this article, we propose a knowledge-embedded contrastive learning framework (KnowCL) for supervised, unsupervised, and semi-supervised HSI classification. The framework diagram is shown in Fig. 2. First, we adopt a disjoint approach to divide the HSIs into training and test sets, and this sampling strategy more realistically reproduces the labeling of HSIs. Then, we obtain a diversity of labeled and unlabeled samples through a novel HSI processing pipeline. The supervised classification model is constructed with an iteration optimization of a backbone and a supervised head. The unsupervised model is optimized to maximize the similarity of the two feature extraction branches. We introduced task-specific knowledge into the unsupervised model to construct a novel semi-supervised learning model, which will greatly speed up the network fitting and enhance the model performance. Moreover, the auto-weighted loss that can adaptively fuse the the supervised learning and contrastive learning is designed for the framework. In addition, we present a light vision transformer (ViT) and test ResNet as backbones to validate the effectiveness of our framework. The main contributions of this research are listed as follows.
  1. 1.
    We combine multiple data augmentation to construct a novel data processing pipeline for HSI classification that well exploits the spectral-spatial information of HSIs and accurately evaluates the performance of the model in realistic scenarios.
  2. 2.
    We propose a new framework for HSI classification named knowledge embedded contrastive learning (KnowCL), which is strongly scalable and can unify supervised, unsupervised, and semi-supervised HSI classification into one end-to-end framework.
  3. 3.
    We develop a new semi-supervised learning paradigm for HSI classification, which is very useful in practical applications. This framework is achieved by a new loss function based on multi-task learning, designed to adaptively combine supervised and contrastive losses.
  4. 4.
    We qualitatively and quantitatively evaluate the classification performance of the proposed model. Extensive experiments indicate that KnowCL achieves performance levels that are on par with or surpass their opponents in supervised and unsupervised methods. Our semi-supervised KnowCL leads all comparison algorithms by a large margin.
The rest of this article is organized as follows. Section II reviews some related works about supervised learning, semi-supervised learning, and unsupervised learning networks. The proposed KnowCL framework is a detailed description in section III. Section IV elaborates qualitatively and quantitatively evaluates the performance of KnowCL on the benchmark datasets. Finally, Section V draws comprehensive conclusions and looks forward to further research directions.
II Related Works
II-A Supervised Learning
In the typical supervised learning classification paradigm, a backbone network is initially employed to extract features from the data. These extracted features are then mapped to a label probability distribution using a softmax layer. The discrepancy between the predicted label and the ground truth is computed by loss functions such as focal loss and cross-entropy loss. The resulting error value is then backpropagated through the network to optimize its parameters. Over the last few years, lots of supervised learning classification methods based on DL have been proposed and become mainstream in HSI classification[29, 30]. Some classic backbones play a huge role in the DL, such as recursion neural network (RNN), CNN and transformer[31, 32, 33].
RNN can only model the spectral features of HSIs, but due to the phenomenon of the same object with different spectra and the same spectra with different objects, it is difficult to achieve good classification results by only spectral features [34], so spatial information was introduced into the HSI classification. For instance, DFFN[35] incorporates spectral-spatial features through global and local feature extraction. The CNN network uses the convolution kernel to scan the image, which can extract the spatial features of the image very efficiently. Especially with the emergence of ResNet [32], the skip connection greatly improves the depth of the CNN and enhances the classification performance. However, CNN can only rely on multiple patches to extract the local spectral-spatial information, leading to a deep and complex network structure. Therefore, FDGC [20] was proposed to extract long-term information in non-European space with GCN, and Hong et. al. [24] presented SpectralFormer with the help of Transformer to conduct HSI classification, which breaks through the limitation of local connectivity in CNN models by self-attention mechanism [36]. They greatly facilitate the development of DL models.
The supervised learning process of DNNs necessitates a substantial volume of labeled training data in order to effectively adjust the model’s parameters. As a result, these models are most suitable for situations where abundant labeled data is available[37].
II-B Unsupervised Learning
The previous unsupervised learning methods [38] extract latent feature representations with the help of image reconstruction. This generation process consumes a lot of resources. As an unsupervised learning method, contrastive learning (CL) has achieved unprecedented performance in computer vision tasks [39, 40, 41]. CL only needs to conduct discriminative tasks in the feature space, thus avoiding the expensive computational cost of generating elaborate fake samples. Therefore, CL models become simpler, and the generalization ability is stronger [42].
Contrastive learning centers on making positive sample pairs closer to each other and negative sample pairs farther apart [43, 44]. For the image, sample pairs can be constructed using image augmentation methods (e.g., rotation, flip, blur, etc.). A sample undergoing a series of data augmentations produces two samples as a pair of positive samples, and other samples from the same batch are viewed as negative sample pairs. MoCo [39] builds a sequence of stored negative samples to expand the number of negative samples. SimCLR [42] uses 128 TPUs directly to train the model to store a large batch size samples. DMVL[45] uses the infoNCE loss function similar to the SimCLR framework to construct the HSI classification model. XDCL[27] learns domain invariance features in the spectral and spatial domains of HSIs.
Unsupervised learning models can directly learn feature representations of HSIs without labels, and it is suitable for large-scale remote sensing image feature extraction. However, unsupervised model training consumes a lot of time, and the extracted task-agnostic features still need labels for specific task fine-tuning.
Figure 3: Data processing overview.
II-C Semi-supervised Learning
While labeled samples in HSIs are currently expensive to obtain, there are a large number of unlabeled samples that are easily accessible. Utilizing a very limited number of labeled samples and a large number of unlabeled samples simultaneously, the semi-supervised method shows good performance [46]. However, how to extract labeled and unlabeled sample features is a problem. There are various solutions, such as constructing pseudo-labels or neural network-based methods [17, 18].
The method based on pseudo-labels can directly use unlabeled samples. It usually clusters HSIs (both labeled and unlabeled) first. After assigning the cluster labels to all unlabeled data, we can train the model using supervised learning [17]. The main problem is that incorrect pseudo-labels cause error accumulation and long training time. Models based on CNNs or GNNs can directly encode unlabeled samples. CNNs [25, 47] extracts features of unlabeled samples through convolution kernels, which is an implicit semi-supervised learning method. By converting HSIs into graph data, features can be directly aggregated on all samples through GNNs [18, 48]. Liu et al. [19] approached HSI classification from a graph perspective, utilizing graph convolutional networks to extract feature representations of irregular image regions in non-Euclidean space. However, these methods necessitate loading the entire HSI into memory, making it unsuitable for large-scale satellite image processing.
III Proposed Approach
The framework incorporates four components, as illustrated in Fig. 2. (a) data processing module: the data processing module prepares the training and test dataset for training and test; (b) encoder module: the encoder module extracts representation vectors from augmented samples; (c) loss module: the loss module generates supervised loss, unsupervised loss, or adaptively fuses supervised and unsupervised loss; (d) predictor module: we can use the predictor module to accomplish the classification task.
III-A Data Processing
We construct a generalized data processing pipeline for real-world HSI classification, which is applicable to supervised learning, unsupervised learning and semi-supervised learning models for HSI classification, including normalization, data splitting, spectral grouping, dimensionality reduction and data augmentation. The structure of this pipeline is illustrated in Fig. 3.
Following the common data processing procedure, we first normalize the data to be between 0 and 1. Then, as stated in Section I, the random sampling strategy is an unrealistic use case. We selected a fixed proportion of HSIs from each class, in turn from top to bottom, as training samples. The remaining labeled samples are used as test samples. The visual division results can be found in IV-A. This method can ensure that the test set information will not be introduced during spatial feature extraction. The ground truth is regional rather than discrete, which allows us to more accurately measure how well the model generalizes to the new land covers.
CL requires learning invariant feature representations of objects from different views of the same object. Data augmentation is widely used for CL as a simple way to generate multi-view representations. HSIs have multiple spectral bands. Beyond conventional data augmentation, these bands can all be viewed as different observations of the same target. We halve the HSI into two groups for extracting invariant feature representation, as shown in Fig. 3. Considering the impact of different sensors and bands on the classification performance, we do not directly reduce each HSI to three dimensions to apply natural image augmentation methods. Instead, we use principal component analysis (PCA) [20] to extract the most discriminative features in each HSI group. With each grouped spectral pixel as the center, image blocks $X\in\mathbb{R}^{H\times{W}\times C}$ are constructed. The common data augmentation includes random crop and resize, random vertical and horizontal flips, and Gaussian blur. It is worth noting that HSIs do not have the concept of color, so data augmentation, such as color jitter, is not adaptive. The input HSI blocks $X$ are divided into two views after data processing: a labeled sample view $x$ and an unlabeled sample view $\hat{x}$.
III-B Encoder Modules
Figure 4: Detailed structures of network components: (a) the attention block, (b) the supervised head, (c) the contrastive head.
TABLE I: Details of the Four Datases.
UP Salinas Dioni DFC2018
Region or Country Pavia California Dioni Houston
Sensor type ROSIS AVIRIS Hyperion ITRES CASI
Bands 103 204 176 48
Wavelength($\mu m$) 0.43-0.86 0.40-2.5 0.4-2.5 0.38-1.05
GSD($m$) 1.3 3.7 30 1.0
Image size 610$\times$340 512$\times$217 250$\times$1376 601$\times$2384
Labeled samples 42776 54129 20024 504712
Classes 9 16 12 20
Release time 2003 2001 2018 2018
The encoder module is composed of a backbone $f_{b}(cdot)$ and two projection heads $f_{s}(cdot)$ and $f_{c}(cdot)$. Our framework allows selecting combinations of network architectures to accomplish supervised classification, unsupervised classification, and semi-supervised classification with an arbitrary network backbone without any restrictions.
Transformer-based methods have shown excellent performance on an increasing number of tasks. Many ViT-based methods have been proposed to solve HSI classification. Especially when the training samples are sufficient, ViT can achieve better performance. Our framework can be trained by global samples, which greatly expands the sample size. Therefore, we choose a simplified ViT as the backbone, named ViT-hsi, which follows the implementation of DINO [41]. Here, we summarize some different network configurations. To encoder images more efficiently, the ViT architecture takes a sequence of flattened 2D patches as input. So we further reshape the image block into 2D patches, which passed through a trainable linear layer to form a set of patch embeddings. Position embeddings and class tokens have also been added to ensure consistency with previous works. Note that the class token only plays the role of aggregating global information when training the supervision head and is no different from other tokens when training the contrastive head. The embeddings and class tokens are passed into a four-layer standard attention block. An attention block is shown in Fig. 4(a). The drop path layer regularization is added after each self-attention layer and multi-layer perceptron (MLP) layer to alleviate over-fitting. The self-attention mechanism updates the feature representation by encoding global image blocks.
CNN-based models also show excellent performance. We also employed several ResNet variants to test the performance of our framework. The sliced HSI block size is usually much smaller than the natural image, so we set all convolutional kernels of ResNet to 3 and removed the pooling layer and the linear layer. Two projection heads $f_{s}(\cdot)$ and $f_{c}(\cdot)$ are shown in Fig 4(b) and (c). The supervised head is responsible for mapping the extracted features to the category space. To improve the accuracy of the mapping, we use a 2-layer MLP. The Dropout layer and Normalization layer is used to improve the generalization ability of the model. The main function of the projection head is to map the features of two contrastive branches into the same subspace for computing the contrastive loss. The projection head consists of an MLP layer with hidden dimension 126 followed by ${l}_{2}$ normalization.
TABLE II: Numbers of Training and Test Samples for UP, Salinas, Dioni, and DFC 2018
UP Salinas Dioni DFC2018
No. Class Train Test Class Train Test Class Train Test Class Train Test
1 Asphalt 1996 4635 Broccoli-weeds-1 419 1590
Dense urban
fabric
632 630 Healthy grass 1458 8341
2 Meadows 5619 13030 Broccoli-weeds-2 806 2920 Mineral extraction site 106 98 Stressed grass 4316 28186
3 Gravel 644 1455 Fallow 425 1551
Non-irrigated
arable land
307 307 Artificial turf 331 353
4 Trees 922 2142
Fallow-rough-
plow
287 1107 Fruit trees 75 75 Evergreen trees 2005 11583
5 Metal Sheets 417 928 Fallow-smooth 540 2138 Olive groves 902 866
Deciduous trees
676 4372
6 Bare Soil 1545 3484 Stubble 794 3165
Coniferous
forest
181 180 Bare earth 1757 2759
7 Bitumen 407 923 Celery 746 2833 Dense vegetation 2540 2495 Water 147 119
8 Bricks 1106 2576 Grapes-untrained 2270 9001 Sparce vegetation 3244 3130
Residential
buildings
3809 35953
9 Shadow 285 662
Soil-senesced-
develop
1311 4892
Sparsely
vegetated areas
895 859
Non-residential
buildings
2789 220895
10 - - - Corn-weeds 684 2594 Rocks and sand 264 228 Roads 3188 42622
11 - - - Lettuce-4wk 230 838 Water 808 804 Sidewalks 2699 31303
12 - - - Lettuce-5wk 414 1513 Coastal water 202 196 Crosswalks 225 1291
13 - - - Lettuce-6wk 185 731 - - -
Major
thoroughfares
5193 41165
14 - - - Lettuce-7wk 228 842 - - - Highways 700 9149
15 - - - Vinyard-untrained 1455 5813 - - - Railways 1224 5713
16 - - - Vinyard-trellis 380 1427 - - - Parking lots1 1179 10296
17 - - - - - - - - -
Parking lots2
127 22
18 - - - - - - - - - Cars 848 5730
19 - - - - - - - - - Trains 493 4872
20 - - - - - - - - - Stadium seats 1313 5511
Ratio(%) - 0.3 0.7 - 0.2 0.8 - 0.5 0.5 - 0.07 0.93
Total - 12941 29835 - 11174 42955 - 10156 9868 - 34477 470235
Figure 5: The training and test ground-truth maps on the four reference datasets, i.e., (a) UP, (b) Dioni, (c) Salinas, (d) DFC2018.
III-C Loss Functions and Fusion Strategy
The existing HSI classification approaches mainly consider cross entropy to assess how close the actual output is to the desired output to calculate the loss value, which is expressed as:
${\cal L}_{ce}(p,q)=-\sum\limits_{x}{p(x)\log q(x)},$ (1)
where $p(x)$ is the expected distribution of $x$ and $q(x)$ is the distribution of the corresponding ground truth. This approach requires ground truth to be involved in each iterative training. In order to take advantage of the rich unlabelled samples, we use CL to extract features of unlabelled samples. We randomly sample a batch size of unlabelled samples to define the CL task. Let $N$ be the size of a batch. The number of unlabelled samples constructed by data augmentation is $2N$. Consider the outputs $z$ and ${\hat{z}}$ of the two branches as a pair of positive samples, then $z$ and the remaining $12N-$ 1 samples constitute negative samples. The loss function is defined as
${\cal L}({z_{i}},{\hat{z}_{j}})=\log\frac{{{e^{\theta({z_{i}},{z_{j}})/\tau}}}}{{\sum\limits_{m=1}^{2N}\mathbb{I}{{{}_{[m\neq i]}}{e^{\theta({z_{i}},{z_{m}})/\tau}}}}},$ (2)
$\theta({z},{\hat{z}})=z^{T}\cdot{\hat{z}}/||z^{T}||\cdot||{\hat{z}}||,$ (3)
where ${z_{i}}$ and ${z_{j}}$ represent the nonlinear embeddings generated from training samples. $\theta({z},{\hat{z}})$ calculate the cosine similarity between $z$ and ${\hat{z}}$. $\tau$ is the temperature parameter and $\mathbb{I}_{[m\neq i]}$ acts as an indicator function. The final optimization goal is to minimize the average over all positive pairs.
${\cal L}_{cl}=-\frac{1}{{2N}}\sum\limits_{i=1}^{N}{[{\cal L}({z_{i}},{\hat{z}_{j}})+{\cal L}({\hat{z}_{j}},{z_{i}})]}.$ (4)
Considering contrastive learning and supervised learning as atomic tasks in a classification task, we can define a multi-task ${\cal T}$, whose loss function ${\cal L}_{T}$ consisting of a cross-entropy loss and an infoNCE loss. This yields a combined loss function
${\cal{L}_{\cal T}}=\sum\limits_{{t}\in{\cal T}}{{\cal{L}}{{}_{t}}\cdot w_{t}}.$ (5)
Due to each of the contributing single loss functions behaving differently, $w_{t}$ is used to regulate the proportion of atomic tasks contributing to the combined loss. Instead of setting a fixed parameter, inspired by [49], we add a learnable parameter to weighting each. This final combined loss function is defined as
${\cal{L}_{\cal T}}=\sum\limits_{\cal{t}\in T}{\frac{1}{{2\cdot w_{t}^{2}}}\cdot{L_{t}}}+\ln(1+w_{t}^{2}),$ (6)
where $\ln(1+w_{t}^{2})$ is a regularization term to avoid trivial solutions.
III-D Training and Test Procedure
As shown in Fig. 2, the training set from the data processing module, encoder module, and loss module is used for training. The supervised loss is computed along the red line. The labeled image block $x$ from the augmented views is passed to the backbone to extract latent features $h$. Then we use a supervised head to obtain predicted labels $p(x)={f_{s}}(h)$. The supervised loss can be calculated by cross-entropy loss between predicted labels and ground truth. The difference between unsupervised and supervised classification is that we use another identical branch to extract features from unlabeled image block $\hat{x}$. Features from both branches are mapped to a high-dimensional subspace using a contrastive head. We maximize their similarity to train the network, as shown by the dark blue line. We design a new semi-supervised learning model relying on fusing supervised and unsupervised learning. During each iteration, we obtain the supervised loss and the contrastive loss separately and then fuse the two in a learnable way to obtain the semi-supervised loss.
After the convergence of KnowCL, the test set is used to evaluate the model performance. A combination of the trained backbone and a classifier is used to validate the effectiveness of the model. Supervised classification directly outputs classification results that can be used for testing. Unsupervised classification usually verifies the model’s effectiveness with the help of fine-tuning. In semi-supervised learning, due to the inclusion of the supervised head, we can directly output the classification results from the supervised head for evaluation in an end-to-end manner. We can also use the linear evaluation protocol [43, 50], which consists in training a linear classifier on top of the frozen representation. Finally, following [51], a simple KNN classifier is used to evaluate the model’s performance. In the experimental session, we discuss each classification approach in detail.
IV Experiments
IV-A Reference Datasets
Four HSI datasets are mainly used in this study which are University of Pavia dataset (UP), Salinas dataset, Dioni dataset and 2018 IEEE GRSS data fusion contest dataset (DFC2018). These four datasets used four sensors and captured four regions at different times, as shown in Fig. 5. The DFC2018 has a minimum of 48 bands, while the Salinas has up to 204 bands. The DFC2018 and UP only scan the visible near-infrared spectra. The Salinas and Dioni acquire both visible near-infrared and shortwave infrared spectra. Dioni is located in the middle of the wilderness and has a low spatial resolution, while the remaining three datasets focus on urban and agricultural scenes and have a high spatial resolution. Of these four datasets, the Salinas was collected the earliest, which is relatively small in size. DFC2018 is approximately 12 times larger than the Salinas data, and also contains the largest number of labeled samples. The detailed parameters associated with these data are given in Table I.
As stated in the section III-A, we adopted a well disjoint data sampling strategy for UP, Salinas, Dioni datasets. For the DFC2018, the training and test set had been defined by the GRSS. The detailed division is shown in Table II and Fig. 5 visualizes the distribution of labels in the training set and test set.
IV-B Implementation Details
IV-B1 Baseline Methods
TABLE III: Quantitative Comparison with Various Crop Sizes on the DFC2018 Datasets.
Crop size 17 18 19 20 21 22 23 24 25
OA 73.63 72.26 73.7 72.15 73.89 73.81 73.98 73.71 73.21
To comprehensively evaluate the performance of the the proposed HSI classification framework in the disjoint data, a traditional machine learning method EMP-SVM and four state-of-the-art models each from supervised learning, semi-supervised learning, and unsupervised learning are selected for comparison, respectively. In the supervised learning scheme, DFFN[35] incorporates spectral-spatial features through global and local feature extraction, and SPRN[26] explores the enhancement of spectral partition on the model performance. FDGC[20] and SpectralFormer[24] explored the application of GCN and transformer modules for HSI classification, respectively. The unsupervised learning AAE and VAE [38] extract latent feature representations with the help of HSI reconstruction for downstream classification tasks. Both DMVL[45] and XDCL[27] use the infoNCE loss function to construct the CL model. For semi-supervised learning, CNN-based FPGA[25] solves the HSI classification problem in terms of semantic segmentation. CEGCN[19], WFCG[18], and EMSGCN[48] are models belonging to GNN, which convert the HSI into a graph by constructing hyperpixels to extract the global features in the non-Euclidean space.
IV-B2 Technical details
This framework unifies supervised learning, unsupervised learning and semi-supervised learning, and we name them KnowCL-SU, KnowCL-US and knowCL, respectively. KnowCL(under kNN classifier) has the best performance and is mainly used to analyse and discuss the superiority of the proposed algorithm. We adopt a modified ViT as the backbone whose has four attention blocks. We first intercept HSI blocks of size $C\times 25\times 25$ based on each center pixel, where $C$ depends on the input HSI. Following the standard ViT, the image blocks are further flattened with patches size of 4. The input HSI blocks will be encoded into 126-D latent features. This features will be further mapped to 256-D latent features by a 1-layer MLP contrastive head and a 2-layer MLP supervisory head to compute the infoNCE loss and cross-entropy loss, respectively. The temperature $\tau$ is set as 0.5. The data augmentation is available in PyTorch’s torchvision package. The random crop size is set to 23×23-pixel, considering the experimental results. AdamW is used as the optimizer with a cosine annealing learning rate schedule. We train for 30 epochs with batch size 256, weight decay 0.000001, and learning rate 0.001. We follow [51, 52] to evaluate the performance of the model with a weighted kNN classifier, and the $k$ value is set to 5 according to the performance of the experiment. We also further tested our model on the supervised head and linear classification protocol to explore the accuracy limits of the model. All models are implemented based on Pytorch 1.9 and performed with an NVIDIA GeForce RTX 3090 GPU.
IV-B3 Metrics
There is no unified model for the HSI classification community, leading to a wide range of performance, number of parameters, and training time for these models. We evaluate these algorithms using common practice in deep learning. They are the overall accuracy, the average accuracy, the Kappa coefficient, parameters, FLOPs, and execution time.
TABLE IV: Quantitative Comparison with Various Batch Size on the DFC2018 Datasets.
Batch size 64 128 256 512 1024
OA 73.78 73.15 73.98 72.91 71.57
TABLE V: Quantitative Comparison with Different Bands on the Four Datasets.
Bands 1 3 5 7 10 15 20 25 30
UP - 69.33 78.23 78.5 91.36 90.11 88.55 85.2 86.52
salina - 87.74 91.96 93.34 92.7 91.12 90.91 91.45 90.93
Dioni 57.44 82.52 80.76 75.79 68.46 68.33 65.01 65.13 63.43
DFC2018 - 64.68 71.675 72.13 73.98 70.85 71.47 - -
IV-C Parameter analysis
TABLE VI: Quantitative Comparison with Other Classification Methods on the UP Dataset.
Learning Parameters FLOPs OA AA Kappa Training time Test time
EMP-SVM - - - 61.58 74.39 53.17 - -
DFFN [35] Supervised Learning 532.7K 175.47M 78.69 79.28 72.64 660.81s 63.03s
SPRN [26] 174.67K 63.43M 78.62 70.54 72.72 270.37s 31.13s
FDGC [20] 2.45M 14.84M 75.63 75.28 68.88 113.98s 2.24s
SpectralFormer [24] 99.25K 10.29M 87.10 87.50 83.24 453.00s 3.23s
KnowCL-SU 931.43K 20.09M 86.27 77.87 81.53 545.99s 3.70
DMVL [45] Unsupervised Learning 11.3M 917.5M 58.75 47.34 45.45 197.67min 7.37s
VAE [38] 824.03K 25.2M 68.40 63.43 59.66 110.53min 43.74s
AAE [38] 758.37K 25.1M 73.64 72.45 66.59 125.1min 28.26s
XDCL [27] 6.47M 77.11M 81.04 82.66 75.38 506.67s 2.17s
KnowCL-US 599.77K 19.76M 82.04 73.23 76.31 516.74 3.16
FPGA [25] Semi-supervised Learning 2.56M 110.42G 63.18 70.73 55.56 83.61s 0.07s
CEGCN [19] 152.85K 13.06G 83.31 87.34 78.83 47.32s 10.10s
WFCG [18] 63.19K 13.23G 79.80 87.38 74.90 60.82s 7.03s
EMSGCN [48] 201.69k 14.68G 86.45 86.25 82.23 64.54s 1.88s
KnowCL(kNN) 997.09K 20.15M 91.36 87.75 88.45 655.76s 3.64s
KnowCL(Linear evaluation) 997.09K 20.15M 91.05 82.99 88.1 33.1s 2.7s
Figure 6: Quantitative comparison with various $k$ values on the four datasets.
(a) UP
(b) Salinas
(c) Dioni
(d) DFC2018
Figure 7: T-SNE isualization of the four datasets using KnowCL.
TABLE VII: Quantitative Comparison with Other Classification Methods on the Salinas Dataset.
Learning Parameters FLOPs OA AA Kappa Training time Test time
EMP-SVM - - - 78.85 84.29 76.28 - -
DFFN [35] Supervised Learning 435.71K 140.31M 82.06 88.95 80.03 473.18s 33.87s
SPRN [26] 176.97K 63.43M 82.06 88.95 80.03 234.56s 51.56s
FDGC [20] 1.98M 14.18M 85.84 91.85 84.21 108.78s 3.92s
SpectralFormer [24] 99.70K 20.29M 87.62 91.90 86.20 551.55s 6.58s
KnowCL-SU 927.19K 19.87M 85.60 90.12 83.93 451.04 3.69
DMVL [45] Unsupervised Learning 11.31M 917.5M 85.06 87.69 83.33 103.47min 5.19s
VAE [38] 824.03K 25.2M 93.07 93.76 92.25 63.27min 38.71s
AAE [38] 758.37K 25.1M 91.15 92.36 90.12 65.30min 45.21s
XDCL [27] 6.67M 99.67M 78.84 82.73 76.54 586.67s 4.89s
KnowCL-US 593.73K 19.54M 93.45 96.34 92.68 447.70 3.68
FPGA [25] Semi-supervised Learning 2.65M 67.77G 78.61 79.76 77.35 81.64s 0.08s
CEGCN [19] 166.89K 8.58G 85.08 88.66 83.26 32.91s 12.19s
WFCG [18] 76.78K 8.62G 85.68 87.84 83.97 26.32s 4.15 s
EMSGCN [48] 215.72K 9.45G 86.92 90.20 85.36 39.37s 1.57s
KnowCL(kNN) 992.85K 19.94M 93.62 96.78 93.11 585.05s 4.47s
KnowCL(Linear evaluation) 992.85K 19.94M 93.53 95.3 92.77 38.3s 3.8s
(a) Ground Truth
(b) SVM
(c) SpectralFormer
(d) EMSGCN
(e) XDCL
(f) KnowCL
(g) False-color
(h) SVM
(i) SpectralFormer
(j) EMSGCN
(k) XDCL
(l) KnowCL
Figure 8: Ground truth (a), false-color (g), labeling area (b)-(f) and full image (h)-(l) classification maps obtained by different models on the UP.
TABLE VIII: Quantitative Comparison with Other Classification Methods on the Dioni Dataset.
Learning Parameters FLOPs OA AA Kappa Training time Test time
EMP-SVM - - - 77.98 69.26 72.97 - -
DFFN [35] Supervised Learning 435.45K 140.31M 57.93 63.70 49.55 438.5s 71.79s
SPRN [26] 176.45K 63.43M 58.62 65.84 50.96 214.11s 46.96s
FDGC [20] 2.45M 14.84M 62.27 75.96 54.16 101.89s 0.84s
SpectralFormer [24] 287.95K 29.55M 74.27 74.69 68.76 389.69s 1.44s
KnowCL-SU 918.09K 19.58M 79.37 78.12 74.73 369.73s 1.66s
DMVL [45] Unsupervised Learning 11.31M 917.5M 60.62 56.78 51.85 368.00min 2.52s
VAE [38] 824.03K 3.23G 58.08 61.49 50.09 141.73min 13.38s
AAE [38] 758.37K 3.22G 66.94 68.39 60.16 170.17min 14.54s
XDCL [27] 6.67M 19.65M 71.05 67.28 64.22 266.67s 1.03s
KnowCL-US 585.66K 19.25M 75.62 72.54 69.95 366.48s 1.21s
KnowCL(kNN) Semi-supervised Learning 983.75K 19.65M 82.52 83.18 78.33 416.86s 1.41s
KnowCL(Linear evaluation) 983.75K 19.65M 81.63 77.22 77.42 17.6s 1.1s
TABLE IX: Quantitative Comparison with Other Classification Methods on the DFC2018 Dataset.
Learning Parameters FLOPs OA AA Kappa Training time Test time
EMP-SVM - - - 45.54 49.91 36.55 - -
DFFN [35] Supervised Learning 435.97K 4.84M 69.04 56.35 60.06 1487.61s 373.67s
SPRN [26] 177.48K 63.43M 66.75 56.29 57.67 786.39s 237.95s
FDGC [20] 2.71M 15.22M 67.74 57.66 58.88 516.34s 42.08s
SpectralFormer [24] 99.97K 4.84M 61.39 67.65 53.61 1870.95s 43.02s
KnowCL-SU 934.26K 20.09M 69.47 54.04 61.58 1351.20s 41.22s
DMVL [45] Unsupervised Learning 11.31M 917.5M 59.16 54.08 49.77 669.11min 482.82s
VAE [38] 824.03K 25.2M 59.55 57.45 50.75 160.25min 5577.63s
AAE [38] 758.38K 25.1M 46.56 53.26 39.09 190.59min 12738.13s
XDCL [27] 6.6M 20.16M 37.92 47.58 30.43 1708.5min 41.6s
KnowCL-US 599.77K 19.76M 62.99 52.37 54.41 1277.37s 36.31s
KnowCL(kNN) Semi-supervised Learning 999.92K 20.16M 73.98 60.6 66.42 2728.18s 43.63s
KnowCL(Linear Evaluation) 999.92K 20.16M 74.08 68.41 66.67 70.21s 45.09s
Deep models have many hyperparameters to be tuned, and these hyperparameters can drastically affect the performance of the model. Conventional hyperparameters such as learning rate, weight decay coefficient, number of hidden layer units, loss function, etc. have been widely and deeply studied [53]. In this study, we focus on four hyperparameters that affect this model. Crop size and batch size affect the overall performance of the model by affecting the feature extraction of the model. The number of bands and k-value of KNN are mainly affected by the dataset. Therefore, we tested the effect of crop size and batch size on the model using the Huston 2018 dataset as a benchmark. The effect of the number of bands and k-value on the model was investigated in each of the four data.
IV-C1 Crop size
Data augmentation affects the quality of unsupervised feature extraction. Chen et al. [42] investigated the effect of different compositions of transformations by the linear evaluation and pointed out that random cropping and random color distortion stand out. Since HSIs do not have the concept of color as in natural images, we only tested the effect of different crop sizes on the HSI classification. Table III shows that the model performs robustly when the images are cropped to different sizes.
IV-C2 Batch size
Table IV shows the impact of batch size when models are trained on the DFC2018. Unlike simCLR [42], model performance does not improve with increasing batch size. On the one hand, it may be due to the fact that the training data involved in the training of the model are fewer, and they are taken from several regions with more similar features so global features can be learned with a smaller batch size. On the other hand, it suggests that the insertion of labeling knowledge allows the model to learn task-specific knowledge, giving the model enough power to find the optimal feature representation under a small batch size.
IV-C3 Bands
HSIs contain bands far beyond those of natural images. The previous work either directly input all the bands or took only 3 bands without considering the effect of different sensors and feature types on the classification performance. However, the spectral correlation affects its classification performance. PCA, a simple dimensionality reduction method is applied to remove the noise interference and extract the most discriminative feature. Table V shows the impact of the number of principal components when the model are trained on the four datasets. We find that when the number of principal components increases, the performance of the model will have a great difference. The overall presentation is an inverse U-shape. The results manifest that different hyperspectral sensors and scenes extract different principal components to achieve the best classification performance.
IV-C4 The value of k
KNN is a simple but powerful non-parametric algorithm, which is often used to evaluate the capability of unsupervised feature extraction. Fig. 6 shows kNN’s evaluation results using different $k$ values for the framework. When different $k$ values are used, similar results are observed on the UP and Salinas. Different $k$ values show opposite trends on the Dinoni and DFC2018 datasets. A slightly large $k$ value will be better on Dinoni, while larger $k$ values are detrimental to DFC2018.
To explore the reason for this, we leverage the t-SNE method to visualize the features extracted by KnowCL for quantitative analysis, as shown in Fig. 7. We can see that due to the purity of the land cover in the first three datasets are more homogeneous and clustered, the features extracted by the model are more separated. Therefore, the model accuracy remains unchanged or improves a little as the $k$ value increases. The DFC2018 has more categories, and the extracted features are superimposed together, resulting in a lower accuracy as more neighboring nodes are taken into account.
IV-D Quantitative Comparison
(a) Ground Truth
(b) SVM
(c) SpectralFormer
(d) EMSGCN
(e) XDCL
(f) KnowCL
(g) False-color
(h) SVM
(i) SpectralFormer
(j) EMSGCN
(k) XDCL
(l) KnowCL
Figure 9: Ground truth (a), false-color (g), labeling area (b)-(f) and full image (h)-(l) classification maps obtained by different models on the Salinas.
(a) Ground Truth
(b) False-color
(c) SVM
(d) SVM
(e) SpectralFormer
(f) SpectralFormer
(g) XDCL
(h) XDCL
(i) KnowCL
(j) KnowCL
Figure 10: Ground truth (a), false-color (b), labeling area (c)-(i) and full image (d)-(j) classification maps obtained by different models on the Dioni.
(a) Ground Truth
(b) False-color
(c) SVM
(d) SVM
(e) SepctralFormer
(f) SepctralFormer
(g) XDCL
(h) XDCL
(i) KnowCL
(j) KnowCL
Figure 11: Ground truth (a), false-color (b), labeling area (c)-(i) and full image (d)-(j) classification maps obtained by different models on the DFC2018.
(a) HSI Patch
(b) ${\cal L}_{cl}$
(c) ${\cal L}_{ce}$
(d) ${\cal L}_{cl}+{\cal L}_{ce}$
(e) ${\cal L}_{\cal T}$
Figure 12: Visualization of different learning strategies output features obtained with ${\cal L}_{cl}$, ${\cal L}_{ce}$, ${\cal L}_{cl}+{\cal L}_{ce}$, and ${\cal L}_{\cal T}$.
The experimental results on UP are shown in Table VI. First of all, we can observe that the difference between the number of model parameters and FLOPs for HSI classification is very large. DMVL uses ResNet50 as the backbone, and its number of parameters is hundreds of times that of WFCG. Traditional semi-supervised learning inputs the whole image into memory, resulting in a dramatic increase in the FLOPs, tens or hundreds of times that of other algorithms. In terms of efficiency, semi-supervised learning is able to perform feature extraction on the whole image simultaneously, so it only takes a few tens of seconds to complete the training. Unsupervised learning, on the other hand, takes a lot of time for each epoch because it trains on every image block. Furthermore, the task-agnostic learning approach also results in poorly learned features for downstream classification tasks.
For supervised and unsupervised learning, KnowCL-SU and KnowCL-US achieve performance levels that are on par with or surpassing comparison methods, while semi-supervised learning KnowCL combines transformer and CL to achieve the best accuracy with balancing time and computation. Moreover, KnowCL is able to extract more fine-grained feature representations than hyperpixel-based graph learning models. Compared to the unsupervised learning model, KnowCL is able to learn task-specific feature representations using label knowledge. In addition, the performance of the linear evaluation is slightly lower than KNN classifier, indicating that our model has fully utilized the label information in the training phase. As shown in Table VII, KnowCL also outperforms the other methods in all quality indices in Salinas. Salinas contains several crops that are uniformly distributed and easily distinguishable, and the other compared algorithms also achieve better classification results.
TABLE X: Comparison under KNN, Supervied Head and Linear Classification Protocols Using Different Backbone Networks on four Datasets.
ViT_hsi ViT_tiny ViT_small ResNet18 ResNet50 ResNet50*2
Parameters 534.11K 5.37M 21.43M 11.17M 23.50M 66.83M
FLOPs 19.69M 198.23M 363.84M 423.00M 984.77M 2.78G
kNN 73.98 74.01 68.55 68.23 67.82 69.65
Supervised head 73.05 74.36 69.69 59.92 55.51 55.5
Linear evaluation 74.19 74.46 70.18 73.35 75.19 78.2
Table VIII and IX show the experimental results on the Dioni and DFC2018. It is worth noting that the memory requirement of traditional semi-supervised learning methods increases exponentially with the amount of data. Conventional GPU are no longer sufficient to support the operation of Dioni and DFC2018. Therefore, only supervised and unsupervised learning methods are used as comparison algorithms. Larger HSIs contain more diverse feature classes and produce more complex spectral combinations, leading to the accuracy of these two datasets having a large gap compared to UP and Salinas. It is in line with the performance of these algorithms when applied to the real world. We can also observe that the linear probe results are higher than the kNN classification results on the DFC2018. This can also be attributed to the complexity of the feature components of DFC2018. The large number of mixed pixels [54] leads to poor classification performance. The best results obtained with KnowCL can substantially outperform comparative models.
IV-E Visualization Comparison
To further evaluate the classification quality of HSIs, we also examined visual perception. Due to space constraints, Figs. 8, 9, 10 and 11 show some of the algorithmic prediction results for the full image and labeled areas on the UP, Salinas, Dioni, and DFC2018, respectively. The classification results obtained by our method are more consistent with the ground truth when comparing Fig. 8(f) and (b)-(e), which demonstrates the effectiveness of the KnowCL. Among the different methods, the edges of the classification map of KnowCL as shown in Fig. 8(l) are smoother than those of the other methods Fig. 8(h)-(k). This indicates that KnowCL can better capture finer spatial detail by the CL and embedded label knowledge. Although EMSGCN obtained clear classification maps by constructing superpixels, as shown in Fig 9[j], this also led to more misclassification of neighboring features, such as Grapes vs. Vineyard and Fallow vs. Stubble. The result in Salinas of KnowCL contains less noise in the classification map, achieving a better visual performance.
The same conclusion can be drawn from the visualisation results of Dioni and DFC2018 in Figs 10 and 11. The Dioni dataset is derived from a spaceborne hyperspectral sensor, and its image quality is more susceptible to atmospheric and ground radiation compared to airborne sensors. Consequently, the classification result maps generated by many algorithms exhibit increased salt-and-pepper noise, as illustrated in Figs. 10(d), (f), (h). KnowCL, however, has the capability to mitigate the influence of local noise signals and produce smooth classification result maps, which is attributed to its training on global data. In the case of the DFC2018 dataset, where 1m resolution remote sensing images are acquired through an airborne hyperspectral sensor, the algorithm faces new challenges due to the complex urban environment. Notably, the limited expressive capacity of XDCL results in a substantial drop in classification accuracy on this dataset. In contrast, KnowCL, equipped with a transformer as the backbone, demonstrates strong expressive ability. The classification result aligns most closely with the genuine feature distribution on the ground.
IV-F Ablation Study
We ablate the loss function to investigate the effect of its on the model. We compared our framework with the ${\cal L}_{cl}+{\cal L}_{ce}$ and ${\cal L}_{\cal T}$ loss function to reveal the effect of the proposed adaptive fusion loss function. As shown in Table XI, when ${\cal L}_{\cal T}$ is used, the final classification results can be further improved in most cases. There is an improvement of 2.48% and 6.09% on the UP and Dioni, respectively. The relatively small boost on the Salinas is perhaps due to the fact that the two branches are quite different and do not complement each other. We observed a slight decrease in the accuracy of the model with adaptive fusion in DFC2018, which is possibly due to the accumulation of errors due to the lower accuracy before fusion.
Fig. 12 visualizes the feature maps using the supervised learning, unsupervised learning, additive fusion learning and adaptive fusion learning strategies, respectively. We selectively pick out some representative feature maps for visual comparison, where the visualization results of using the ${\cal L}_{cl}+{\cal L}_{ce}$ and ${\cal L}_{\cal T}$ learning strategies have finer appearance than those using only one loss function. For example, the activation maps for the region in the white box show that the non-residential buildings area based on the fusion loss has sharper edges, object contours, and texture structures. This also visually confirms the superiority of the designed framework.
TABLE XI: Quantitative Evaluation of Different Learning Strategies on the Four Datasets.
UP Salinas Dioni DFC2018
${\cal L}_{cl}+{\cal L}_{ce}$ 88.88 93.42 76.43 75.94
${\cal L}_{\cal T}$ 91.36 93.62 82.52 73.98
IV-G Backbones and classifier
There are numerous models for HSI classification at present. We try to propose a general framework which is compatible with the mainstream deep learning backbone. Table X shows the performance of three ViT-based networks and three ResNet-based networks as well as their number of parameters under kNN, supervised head and linear classification protocols. Fig. 1 provides a more visual comparison of them with the other models in DFC2018. We find that the ViT-based networks have similar classification performance under the three classification strategies, while the accuracy of the ResNet-based networks becomes higher at linear evaluation as the model parameters increase. We surmise that this is because ViT is able to extract global features at each layer, so the model performance converges faster. ResNet has the local receptive fields and requires deeper structure to adequately extract global features. Moreover, our model achieves a large advantage in all parametric quantities.
V Conclusion
HSI classification techniques have been developed over a long period. A range of models have now been developed. However, the quality of labels and the learning approach result in low accuracy of supervised and unsupervised learning methods in the real world. Traditional semi-supervised learning methods can utilize both labeled and unlabeled samples, but they lack the scalability to perform large-scale HSI classification tasks.
In this paper, we presents a universal HSI classification framework designed for addressing real-world HSI classification problems, which is compatible with supervised learning, unsupervised learning and semi-supervised learning. Among them, semi-supervised learning classification KnowCL is newly defined to fit the requirements of real applications. An efficient contrastive learning strategy is constructed by means of labeled knowledge embedding, where the parameters, FLOPs and execution time is comparable to that of supervised learning but can fully extract features from unlabeled samples. We conduct four disjoint datasets, and a large number of experiments on them quantitatively and intuitively validate the superiority of the proposed model. Future, it is possible to adopt pre-trained visual foundation model for HSI classification, which will greatly advance the application of HSIs in the real world.